Aldean Corporation has shares of stock trading at $51.09 today. The company has 3.00 million shares outstanding. Earnings for the company last year were $7.00 million. Construct a P/E multiple for the company.
Answer #1

P/E=(Price per share)/(Earnings per share)
Earnings per share=(7 million)/(3 million)=7/3
Price=$51.09

P/E=51.09/(7/3)=21.8957 or 21.90 (Rounded to two decimal places)

So, the PE multiple is 21.90